40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 141 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  9992   Mon May 26 07:59:23 2014 KojiUpdateElectronicsAmplifier removed from BeatX path

And the out-of-loop level of the ALSX compared with the previous measurement is ...?

Quote:

I just realized that I forgot to elog this, but yesterday afternoon I bypassed the amplifier in the BeatX path, and now the X beatnote is about -27dBm.  Arms lock nicely with ALS.

 

  9991   Sat May 24 22:56:57 2014 JenneUpdateElectronicsAmplifier removed from BeatX path

I just realized that I forgot to elog this, but yesterday afternoon I bypassed the amplifier in the BeatX path, and now the X beatnote is about -27dBm.  Arms lock nicely with ALS.

  9990   Fri May 23 11:58:28 2014 manasaUpdateGreen LockingY arm green alignment tuned

The Y arm green transmission had come down to 0.3 and the green steering mirrors on the Y end table required some minor alignment adjustments to bring back transmission to around 0.75 counts.

  9989   Thu May 22 11:21:06 2014 SteveUpdateSUSETMX oplev

Quote:

Quote:

Anodized aluminum dumps replaced by 6 razor beam dumps.

Two more razor beam dumps added this afternoon.   The picture will updated tomorrow.

 There are 9 razor beam dumps at ETMY-ISCT

 I added two green glass absorbers. The oplev centering may need a touch up when it is well aligned.

Attachment 1: ETMXoplev.png
ETMXoplev.png
  9988   Thu May 22 00:30:40 2014 manasaUpdateLSCALS X noise from angular motion of mirrors

Below are the transfer functions measured between the angular (pit, yaw) motion of X arm mirrors and the ALSX error signal. The measurements were again made for 1Hz-30Hz.

The transfer functions are mostly flat.

ITMX P - 200-300 Hz/urad (beat freq = 45 MHz)

ITMX Y - 700-800 Hz/urad (beat freq = 27MHz)

ETMX P - 500-600 Hz/urad (beat freq = 56 MHz)

ETMX Y - 1000-2000 Hz/urad (beat freq = 62.5MHz)

Data xml files can be found in /users/manasa/data/140521/

Attachment 1: ALSX_OLPerrITM2.png
ALSX_OLPerrITM2.png
Attachment 2: ALSX_OLYerrITM.png
ALSX_OLYerrITM.png
Attachment 3: ALSX_OLPerrETM.png
ALSX_OLPerrETM.png
Attachment 4: ALSX_OLYerrETM2.png
ALSX_OLYerrETM2.png
  9986   Wed May 21 22:15:37 2014 ericqUpdatePSLPMC relocked

PMC has been unlocked for ~4hrs, not sure why. It's servo gain was down at -10dB...

Relocked with transmission of .76V, MC locks fine with WFS, transmission of 15.5k.

  9984   Wed May 21 13:51:19 2014 SteveUpdatesafetyscope battery recall

Tektronix RECALL on TDS3000 or TDS300B  oscilloscope BATTERIES TDS3BATB

This Lithium-Ion battery can be a fire hazard !  Remove battery pack and recycle it through Safety Office !

  9983   Wed May 21 13:20:34 2014 manasaUpdateLSCALS X noise from angular motion of mirrors

Quote:

[Rana, Jenne]

9.  Arm cavity alignment.  Significant DC effect.

    * When the alignment of one of the arm cavity mirrors is changed, the DC value of the beatnote signal changes. 

           * ITMX moved in yaw, we see a 7kHz/15urad DC shift in the BEATX_FINE_PHASE_OUT_HZ time series.

          * ETMX moved in yaw, we see an 8kHz/5.5urad DC shift in the time series.  We aren't sure why this is about a factor of 3 times larger effect (same shift for smaller misalignment) than the ITM.

    * We want to do a Yuta-style analysis to see what the angle to length coupling looks like, so that we can measure the angular motion of our cavity mirrors and put the expected noise into our ALS noise budget.  Perhaps this will help us understand the low frequency difference between our in-loop beatnote error signal and our in-loop PDH error signals (red vs. maroon on the ALS noise budget posted above Pianosa). 

    * I've asked Manasa to take some transfer functions in the morning, so that we can start to have an idea of what is going on with this.

Attached is the measurement of the transfer function from ITMX oplev error in yaw to the ALSX error signal.

The arm was locked to the IR using POX and the green beat frequency (between X arm trans in green and PSL green) in this case was 27MHz.

The transfer function looks mostly flat between 1Hz - 30Hz at 700-800 Hz/urad. The DC shift that Jenne measured from the time series is ~500 Hz/urad.

So far I have not been able to measure the TF below 1Hz without the arm losing its lock. Updates will follow.

Data xml file can be found in /users/manasa/data/140521/

Attachment 1: ALSX_OLYerrITM.png
ALSX_OLYerrITM.png
  9982   Wed May 21 13:18:47 2014 ericqUpdateCDSSuspension MEDM Bug

I fixed a bug in the SUS_SINGLE screen, where the total YAW output was incorrectly displayed (TO_COIL_3_1 instead of TO_COIL_1_3). I noticed this by seeing that the yaw bias slider had no effect on the number that claimed to be the yaw sum. The first time I did this, I accidently changed the screen size a bit which smushed things together, but that's fixed now.

I committed it to the svn, along with some uncommitted changed to the oplev servo screen.

  9981   Wed May 21 11:11:01 2014 manasaUpdateIOOMC tuned

I found MC unlocked this morning. I looked at the 2 day trend of the MC suspensions and found MC2 suspensions have been misaligned.

I used Rana's ezcaservo trick to recover MC lock. This brought the MC_REFL down to 0.7 counts. I did the rest of the alignment by moving the MC2 suspension sliders only. MC_REFL is down to 0.45-0.5 counts and TRANS_SUM is at ~16300.

Also, I found the WFS servo was left turned OFF. I re-enabled them as well.

Attachment 1: MC_SUS2days.png
MC_SUS2days.png
  9980   Wed May 21 09:35:33 2014 SteveUpdateSUS razor beam dumps ETMX_ISCT

Quote:

Anodized aluminum dumps replaced by 6 razor beam dumps.

Two more razor beam dumps added this afternoon.   The picture will updated tomorrow.

 There are 9 razor beam dumps at ETMY-ISCT

Attachment 1: 9rbdETMY.jpg
9rbdETMY.jpg
  9979   Wed May 21 05:05:39 2014 JenneUpdateLSCREFL 165 vs 33 investigations

[Rana, Jenne]

We spent some time tonight looking at locking the PRMI with REFL165 vs. REFL33, while reducing the CARM offset. 

We were not able to lock the PRMI on REFL165 I&Q at small CARM offsets.  When locking at larger CARM offsets (about 100 counts, which is about 100nm) and then re-adjusting the REFL165 demod phase as I reduced the CARM offset, I saw that I had to significantly rotate the phase.  For PRMI only (no arms), the REFL165 demod phase was -138.5 deg.  When the PRMI was locked with a -100 count CARM offset, the optimal demod phase was -123 deg.  Then at -90 counts the phase was -113 deg.  At -70 counts, the phase was -108 deg, at -50 counts it was -98 deg, and at -40 it was -93 deg.  We want to go back and look at these more carefully, and in a more continuous way, by watching the sensing matrix calibration lines.  It's unclear to me right now why we're seeing this, but it's possible that we're getting some kind of extra 55MHz resonances.

REFL DC looks like it should be good - same slope and gain as sqrtTR, extra 20 or 30 deg of phase margin, so we think that we should be able to transition over to it, and then try engaging the AO path.  Tonight we had Den's new 1kHz lowpass engaged, and with this, everything looks nice and stable.

Game plan:  Bring CARM in until transmissions are at about 10ish, then try keeping CARM on sqrtInvTrans for the DC part, and engage the AC AO part with REFL DC.  We probably just need to try this for a while more to find just the right way to turn it on.

Need to think about demod phase rotation vs CARM offset as well as extra resonances, but this may take a while, and if we can just get the AO path engaged, that would be good.

  9978   Wed May 21 00:18:40 2014 JenneUpdateLSCPRMI locked with REFL33 vs. REFL165

Since Q has found that REFL165 will be better for holding the PRMI while we reduce the CARM offset, I had a look at locking PRMI sideband locking with both 3f PDs.

I checked the REFL165 demod phase, and changed it from -142.5 deg to -138.5 deg. to minimize the Q signal while driving PRM length.

I found that keeping the MICH and PRCL loop gains the same, and using matrix elements +0.1 for both I and Q for REFL165, rather than +1 for both I and Q for REFL 33.

MICH gain is +0.8, PRCL gain is -0.02.  FMs 4,5 on for both, FM 2 triggered for MICH, FMs 2,3,6 triggered for PRCL.

I then locked the PRMI on sideband with REFL 33 and then REFL 165, and measured the other one as an out of loop sensor of the motion.  I find that REFL33 and 165 are both comparable, and so we shouldn't have any trouble using REFL165 for locking.

PRMI_outofloop_20May2014.pdf

  9977   Tue May 20 22:42:28 2014 ericqUpdateLSC3f Stability

 Here's the angles of MICH and PRCL from the my earlier plot by themselves; this shows that the individual demod angles in REFL165 aren't changing much either. 

PRCangles.pdf

  9976   Tue May 20 16:48:52 2014 ericqUpdateLSC3f Stability

So, I really should have done this as soon as Manasa measured the arm lengths... I've updated my MIST model with the real arm lengths, but still am using assumed identical losses of 75ppm on each mirror. (I've tried measuring the arm losses for real, but got numbers in the hundreds of ppms, so I need to reexamine things...) 

Here's a simulation of the fields in a perfectly locked PRC when CARM is swept (Normalized to input power = 1). 

CARMsweepPrcFields.pdf

More importantly, here's the latest simulation of MICH vs. PRCL demodulation angle separation in the 3F signals. It seems that we may be getting burned by using REFL33 for the PRC lock. REFL165, on the other hand looks much more robust. We should try this out. 

3fs.pdf

(Some of my previous simulations incorrectly implemented MICH excitations; I only moved the ITMS, not the ETMS along with them, so some other stuff slipped in... )

  9975   Tue May 20 15:54:39 2014 ranaSummarySUSETMX oplev qpd board

 This QPD circuit (D980325-C1 ) uses the nice OP497 Quad FET opamp as the transimpedance amplifier. It has a low enough current noise, such that we can increase the resistors (R1-4) up to 100k and still be Johnson noise limited. We should also make sure that the compensation caps (C3-6) are ~2.2 nF so as to not destabilize the opamp. f_low = 1/2/pi/R/C = 730 Hz.

I will do the swap later today unless someone else gets to it first. (note: check for oscillations w/ fast scope probe after installing)


I did these modification tonight. The slideshow of some images is attached. Instead of 100k, I used 97.6k thin film, since this seemed like an oddball size that doesn't get used otherwise. I forgot to measure the dark noise of the quadrants before doing the swap, but comparing the pit/yaw/sum before/after the swap shows that the signal is basically unchanged (since pit/yaw is normalized by SUM), but that the noise is lower by a factor of a few above 100 Hz due to being above ADC noise now. Previously, it was bottoming out at ~10 prad/rHz. Since the signal is unchanged, I guess that the calibration and therefore the loop gain should not have changed either...

And the sum went up by almost 10x as expected from the resistor change.

  9974   Tue May 20 11:48:22 2014 SteveUpdateSUSbeam dumps added to ETMX_ISCT

Anodized aluminum dumps replaced by 6 razor beam dumps.

Two more razor beam dumps added this afternoon.   The picture will updated tomorrow.

Attachment 1: razorbeamdumps.jpg
razorbeamdumps.jpg
Attachment 2: razorbeamdumpss.jpg
razorbeamdumpss.jpg
  9973   Tue May 20 09:12:25 2014 SteveSummarySUSETMX oplev qpd board

Quote:

 ETMX oplev qpd gain has to be increased.

 

 Atm3, Oplev sum read 12,000 counts when the qpd was disconnected ?

           Dark qpd was zero and normal He/Ne incident on qpd was 1,730 counts.

Attachment 1: ETMXOLQPD121.jpg
ETMXOLQPD121.jpg
Attachment 2: ETMXOLQPDface121.jpg
ETMXOLQPDface121.jpg
Attachment 3: ETMXolSUM.png
ETMXolSUM.png
  9972   Tue May 20 02:12:35 2014 JenneUpdateLSCALS X noise investigation

[Rana, Jenne]

We have looked at a few things that do and don't affect the out of loop noise of the ALS X beat, and found that cavity alignment and beatnote RF frequency had the strongest effects.

Possible causes of noise:

1.  Air currents from A/C or flowbench.  No effect

        * When table lid is on, turning on and off the flow bench air did not qualitatively change the out of loop beatnote time series signal.

2.  Scattered light from other beams hitting green PDH PD.  No effect.

        * There are a few spots of green light that are hitting the case of the PDH photodiode, but when I put an iris in place to block those spots, there was no change in the beatnote spectra.  This makes sense to me since none of those spots were close to hitting the diode itself. 

        * Rana did notice that the beam was not well centered on the PD, so he steered the beam onto the center of the diode.  Also, the PD is now tilted a little bit so that the reflection from the diode doesn't go back into the beam path.  Neither of these things had an effect that we noticed in the beatnote noise.

3.  Oplev laser light getting to PDH PD.  Not tested.

       * We don't see any red light over by the PDH PD, so we did not try turning off the oplev's laser to see if that had an effect, but we suspect that it is not the cause of our noise.

4.  Clipping of main IR / green beam on Xend table.  Not tested.

      * We should still go have a look at this, but we no longer think that this is the main cause of the elevated noise.

5.  Scattered light all over Xend table.  Not tested.

     * We should still work on dumping extraneous beams on the table, but we do not think that this is the main cause of the elevated noise.

     * Rana took some photos so that we can see how truely bad the situation is.

6.  Amplitude modulation dip in NPRO.  Not tested.

    * It is probably still a good idea to check this, in case the dip in the amplitude modulation has changed over the year or two since it was last measured, but we also don't think that this is the main problem.

7.  Check PDH servo.  Not done.

     * I think this is still on Q's long-term todo list, but we should give the PDH servos a once-over.

8.  Arm cavity longitudinal motion.  No effect.

     * While the Xarm was locked with IR, we put a line at 1.7 Hz with 325 counts into the ITMX position.  To keep lock, the ETM had to move as well.  When we turned on this line (and increased its amplitude up to the final value of 325 cts), we did not see any qualitative change in the beatnote time series noise.

9.  Arm cavity alignment.  Significant DC effect.

    * When the alignment of one of the arm cavity mirrors is changed, the DC value of the beatnote signal changes. 

           * ITMX moved in yaw, we see a 7kHz/15urad DC shift in the BEATX_FINE_PHASE_OUT_HZ time series.

          * ETMX moved in yaw, we see an 8kHz/5.5urad DC shift in the time series.  We aren't sure why this is about a factor of 3 times larger effect (same shift for smaller misalignment) than the ITM.

    * We want to do a Yuta-style analysis to see what the angle to length coupling looks like, so that we can measure the angular motion of our cavity mirrors and put the expected noise into our ALS noise budget.  Perhaps this will help us understand the low frequency difference between our in-loop beatnote error signal and our in-loop PDH error signals (red vs. maroon on the ALS noise budget posted above Pianosa). 

    * I've asked Manasa to take some transfer functions in the morning, so that we can start to have an idea of what is going on with this.

10.  Beatnote RF frequency.  Significant broadband effect.

     * We have found that when the Xarm beatnote is at low RF frequencies, the noise is high, and when the beatnote is at high RF frequencies, the noise is low! 

     * Low RF freqs are below about 40 MHz, while high RF freqs are above about 90 MHz.  This has not been tested for the Yarm.  Also, these are for the case of "temp slider up, beatnote up".  I have not checked if the same is true for the other side of the PSL frequency, although I don't have reason to believe that it would be.

     * Maybe we are saturating some amplifiers?  We need to check this out.  One thought that Den mentioned was the harmonics, and that perhaps they are causing trouble in the electronics.

     * Den is going to think about implementing a frequency divider so that we can directly digitize the beatnote signal. 

    * Here are spectra for different cases:

          ALS_outofloop_19May2013.pdf

        * And here is a spectrogram showing us going back and forth between the high and low noise states:

          XbeatSaturate.png

                     *  A:  First noticing that noise is good when RF frequency is high.

                     * B:  Not locked on TEM00 mode, so extra noisy.  Disregard.

                     * C:  Bad noise time.  Xbeat was 21 MHz (dark purple on DTT spectrum above), Ybeat was 118 MHz (sea green on DTT spectrum above).

                    * D:  Good noise time. Xbeat was 89 MHz (light purple on DTT plot), Ybeat was still 118MHz (turquoise on DTT plot).

                     * E:  Bad noise time.  Xbeat was 37.5 MHz, Ybeat was still 118 MHz.

                     * F:  Good noise time.  Xbeat was 113 MHz, Ybeat was still 118 MHz.

  9971   Mon May 19 22:44:21 2014 denUpdatePEMxend sei isolation kit is set

Quote:

Yend seismometer isolation kit (elog 8461) hosts Guralp seismometer. I made a cable for inside connection, assembled the kit and relocated the instrument from its previous position at the yend inside the kit.

Seismometer is connected to the readout box and running.

    

 Xend internal cabling and external connector is ready. We are waiting for seismometer from Gyro lab. We still need to fix the pot with clamps after we put the instrument in.

We also need a long cable from Xend to the guralp readout box.

  9970   Mon May 19 09:19:44 2014 SteveUpdateLSChappy IFO

15 hours

Attachment 1: whenthinksareworking.png
whenthinksareworking.png
  9969   Sun May 18 17:57:54 2014 ranaUpdateLSCETMX transients due to unseated bias cable

 The daytime crew had noticed that there were some ETMX angular shifts happening without any control or intention.

I reseated and strain relieved the bias cable coming into the backplane of the coil driver and now it seems OK.

In the 4-hour-long second trend plot below, the era before 2300 is before reseating. Afterwards, we make a couple adjustments, but so far there has been no un-asked for alignment shifts.

AdS has been run on both arms and offsets saved. Its locked on green/red and the beat frequencies are low and the amplitudes high. 

Sun May 18 23:53:37 2014: still OK...I declare it fixed.

 

Attachment 1: ETMXbiasCableSeated.pdf
ETMXbiasCableSeated.pdf
  9968   Sun May 18 14:36:04 2014 denUpdateLSCoffsets in 3f and drmi stable lock on 1f

Eric, Den

We noticed that PRMI RIN is high when the cavity is locked on RELF33 I&Q signals. We compared the level of power fluctuations when PRMI was locked using REFL11, REFL55 and REFL 33. Attached plot "prmi_rin" shows the spectra.

Then we excited PRM and measured length to RIN coupling when PRMI was locked on REFL33 I&Q. DC offset of PRCL is only 3%. But MICH offset seems to be ~nm. When we gave offset of -15 cnts to the servo, power fluctuations improved by a factor of few.

Then we looked at DRMI. It seems that SRC macroscopic length is off but we still could lock it stably. To account for macroscopic length detuning we had to rotate REFL55 phase from 25 degrees to 50 degrees. Power at AS port increased by factor of ~2 compared to PRMI configuration. SPOP18 is decreased only by 30%. Attached plot "drmi_power" shows POP18, POP90, POPDC and ASDC in PRMI and DRMI configurations.

Plot "src_ol" shows srcl OL transfer function. UGF is 70 Hz. We have also centered SRM OL and copied the servo filters from PRM, gains are set to keep UGF at ~0.1 Hz and 7 Hz

This is a more detailed procedure:

1. Phase: REFL11: 19 degree, REFL55: 50 degrees (25 degrees for PRMI configuration)

2. Input matrix:

PRCL   0.15 0 0 REFL11I
MICH = 0 0.15 0 REFL55Q
SRCL   -0.09 0 1 REFL55I

3. Servo parameters:

PRCL: gain = -0.02, FM4,5 + trigger FM2,3,6,9

MICH: gain = 0.8, FM4,5 + trigger FM2

SRCL: gain = 0.25, FM4,5 + trigger FM2,3

4. Triggering:

signal is SPOP22 , levels 40:25

Attachment 1: prmi_rin.pdf
prmi_rin.pdf
Attachment 2: drmi_power.png
drmi_power.png
Attachment 3: src_ol.pdf
src_ol.pdf
  9967   Sat May 17 14:48:06 2014 denUpdatePEMyend sei isolation kit is set

Yend seismometer isolation kit (elog 8461) hosts Guralp seismometer. I made a cable for inside connection, assembled the kit and relocated the instrument from its previous position at the yend inside the kit.

Seismometer is connected to the readout box and running.

IMG_1405.JPG    IMG_1406.JPG

  9966   Fri May 16 20:55:18 2014 JamieFrogsloreun-full-screening Ubuntu windows with F11

Last week Rana and I struggled to figure out how to un-full-screen windows on the Ubuntu workstations that appeared to be stuck in some sort of full screen mode such that the "Titlebar" was not on the screen.  Nothing seemed to work.  We were in despair.

Well, there is now hope: it appears that this really is a "fullscreen" mode that can be activated by hitting F11.  It can therefore easily be undone by hitting F11 again.

  9965   Fri May 16 16:08:12 2014 steveUpdateLSCIsolating base plates

 Electronic components should be ISOLATED as they are installed on the optical tables.

This is essential to avoid ground loops, 60 Hz and harmonic peaks in the spectrum. We have just got some made.

Please only use it for this reason. Earlier black delrin base plates were used up in not needed places.

 

The anodized Aluminum base plates with magenets certainly will conduct.

 

Attachment 1: ISObaseplates.jpg
ISObaseplates.jpg
  9964   Fri May 16 11:22:23 2014 SteveUpdateLSCX Arm ALS Noise coming in and out

Quote:

Den and I spent some time with the interferometer last night with hopes of bringing in the AO path, but were stymied by the (re)occurrence of the anomalously high low frequency motion of the Xarm, as seen by fluctuations of TRX from .9 to .2 while "held" on resonance.

Jenne reported that they weren't seeing it earlier in the evening, and then it started again when I showed up. Holding the arms on IR, we could see a fair amount of excess low frequency noise in the BEATX_FINE_PHASE_OUT_HZ channel, as compared to BEATY, bringing its RMS to 5 times that of the Y arm. From the shape of the excess noise (broad slope from DC to tens of Hz), Rana suspected air currents and/or scattering effects being the culprit.

Den poked around a bit on the PSL table, which didn't really change much. He then went down to the X end table to inspect the table, and while he was there, I noticed the noise go down to being in line with the Yarm. I joined him at the end, and we found the beat phase noise in the frequency region of concern to be hugely sensitive to tapping on the enclosure, air current, etc. There is also a ton of green light everywhere, and multiple spots of green light around the green refl PD.

At that point however, the quiescent noise was acceptable (TRX fluctuations of <.2), so we went back to the control room to try to lock. Unfortunately, after a few attempts, the noise was back. At this point, we went home. The layout of the end table likely needs some attention to try and minimize our susceptibility to excess scatter effects. 

 Turn off the AC and flow bench please.

  9963   Fri May 16 10:54:42 2014 ericqUpdateLSCX Arm ALS Noise coming in and out

Den and I spent some time with the interferometer last night with hopes of bringing in the AO path, but were stymied by the (re)occurrence of the anomalously high low frequency motion of the Xarm, as seen by fluctuations of TRX from .9 to .2 while "held" on resonance.

Jenne reported that they weren't seeing it earlier in the evening, and then it started again when I showed up. Holding the arms on IR, we could see a fair amount of excess low frequency noise in the BEATX_FINE_PHASE_OUT_HZ channel, as compared to BEATY, bringing its RMS to 5 times that of the Y arm. From the shape of the excess noise (broad slope from DC to tens of Hz), Rana suspected air currents and/or scattering effects being the culprit.

Den poked around a bit on the PSL table, which didn't really change much. He then went down to the X end table to inspect the table, and while he was there, I noticed the noise go down to being in line with the Yarm. I joined him at the end, and we found the beat phase noise in the frequency region of concern to be hugely sensitive to tapping on the enclosure, air current, etc. There is also a ton of green light everywhere, and multiple spots of green light around the green refl PD.

At that point however, the quiescent noise was acceptable (TRX fluctuations of <.2), so we went back to the control room to try to lock. Unfortunately, after a few attempts, the noise was back. At this point, we went home. The layout of the end table likely needs some attention to try and minimize our susceptibility to excess scatter effects. 

  9962   Fri May 16 10:48:32 2014 SteveUpdateLSCY arm T- qpd is getting light guard extension tube

The qpd was removed from the east end table and threaded adaptor ring epoxied on it's shell.

This tube will cut down the amount of emergency exit light getting into the qpd.

Attachment 1: TRYepox.jpg
TRYepox.jpg
  9961   Fri May 16 09:46:05 2014 SteveUpdatePSLpointing monitoring

Quote:

 Tonight I noticed that the drop in PMC transmission was ~1V, more than the usual of ~0.5V from the daily drift.

While re-aligning on the table, I noticed that the misalignment was not from either of the steering mirrors; i.e. I has to walk them both to get the alignment back. This implies that the misalignment is generated far upstream. Maybe the the laser itself is moving. We need some updates from Steve's laser misalignment tracker.

I'd like to replace the paper target with IOO -QPD_POS so we can log it.

  9960   Fri May 16 00:25:53 2014 ranaUpdatePSLPMC realign

 Tonight I noticed that the drop in PMC transmission was ~1V, more than the usual of ~0.5V from the daily drift.

While re-aligning on the table, I noticed that the misalignment was not from either of the steering mirrors; i.e. I has to walk them both to get the alignment back. This implies that the misalignment is generated far upstream. Maybe the the laser itself is moving. We need some updates from Steve's laser misalignment tracker.

  9959   Thu May 15 16:46:35 2014 ericqUpdateLSCPossible Path to AO path

 It's taken a lot of trial and error, but I've found a path through MATLAB loops that seems like it may be stable at all points.

CAVEAT: This doesn't give any indication as to why we weren't able to turn up the AO gain more last night, as far as I can tell, so it's not all good news. 

However, it's still ok to at least have a plan that works in simulation... 

Based on the location of the optical resonance peak in the CARM plant, we estimate our CARM offset to be 200pm. I haven't simulated TFs there exactly, but do have 100pm and 300pm TFs. This procedure works in MATLAB starting at either, though 100pm is a little nicer than 300. MATLAB data and code is attached in a zip. 

The steps below correspond to the attached figures: Bode plots and step response of the Loop at each step. 


0. [Not Plotted] DCTrans sensing, MCL actuation on CARM. FMs1,2,3,5,8; UGF = 120. (DARM not considered at all)

  1. AO path just turned on. Crossover with MCL path ~ 3.5kHz. 
  2. AO gain increased. Crossover ~ 500 Hz. There are now multiple UGFs! Handling all of these in a stable manner is tricky. 
  3. AO gain increased. Crossover = 150Hz. [No simulations with a higher crossover survived the next steps]
  4. Compensation filter applied to MCL path; 1 real Zero at 105Hz and a pole at 1k. From a TF point of view, this is sort of like switching to REFLDC, but the SNR at low frequencies is probably better in TR signals at this point. 
  5. CARM offset reduced to 30pm. (This smoothens out the optical plant resonance.)
  6. Overall gain increased by factor of 3. There is now just one UGF at a few kHz, above the optical resonance. From here, gain can be further increased, boosts can go on, offset can go way down. In reality, we should switch to a single error signal once we're back to one UGF, and go from there. 

AOtransitionBodes.pdfAOtransitionSteps.pdf


#4 Seems like the most sticky part. While both sides of this look stable as far as I can tell. I feel that flipping from the red phase curve to the teal might not actually be ok, since they are on either side of the bad phase of 0 degrees. It isn't immediately evident to me how to easily model the transitions between steps, rather than just the stability of of each step in the steady state. 

Attachment 3: AOCarmLoop.zip
  9958   Thu May 15 15:10:02 2014 SteveUpdatePEMseismic activities

Our only seismometer is at the east end now.

Atm1, Ditch Day morning puzzle. The gardener came after the freshman did leave and cut the grass with the lawn mower.

Atm2, Yesterday afternoon the Aztecs containers moved out.

Attachment 1: DichDay.png
DichDay.png
Attachment 2: Ditchday8am.jpg
Ditchday8am.jpg
Attachment 3: AztecsGone.jpg
AztecsGone.jpg
  9957   Thu May 15 02:52:51 2014 JenneUpdateLSCStarted engaging the AO path, not getting all the way yet

 

 In addition to a transparent legend, we need the corresponding CM crossover measurements from DTT to compare with the Q-Mist-Loop model results. The xover tells us when the AO gain is high enough so that they can be ramped up together.

Also, I wonder how much power fluctuation we get from the large ALS DIFF noise and if that demands we get the TR signals normalized by POPDC.

  9956   Thu May 15 02:32:01 2014 JenneUpdateLSCStarted engaging the AO path, not getting all the way yet

I tried many times this evening to engage the AO path, with limited success.

Q's new scripts worked really well, and so I have some transfer functions!  To take these measurements, in ...../scripts/general/netgpib, I am running ./TFSR785 TFSR785_CARMloop_May2014.yml, where the file name is the name of my parameter file.  The data, and the saved pdfs, are in /users/jenne/PRFPMI/CARM_loop_measurements/2014May14/ .  For these measurements, the SR785 is hooked up to the "A" set of excitation and test points on the CM board.

All of these traces were taken while the IFO was PRFPMI, with PRCL and MICH on REFL33, DARM on ALS diff, and CARM on InvSqrtTrans.  carm_cm_up.sh is up to date, through the echo "REFL_I should now be zero" (~line 111 in the script).  All you need to do is set the beatnotes, and then run the script.  Follow instructions in the prompt (such as "press enter to confirm PRMI is locked"). 

TFSR785_15-05-2014_011008.pdf

Here are my notes for the various times:

23:01:44 - MC IN2 = 0dB, CARM gain = 5.0

23:13:45 - MC IN2 = 10dB, CARM gain = 5

23:26:10 - MC IN2 = 6dB, CARM gain = 10 (after Q suggested increasing overall gain, rather than just AO path)

00:13:07 - MC IN2 = 6dB?, CARM = 6ish?  don't remember exactly.

00:45:00ish, Realigned IFO using IR with arms.

01:03:17 - MC In2 = 0dB, CARM gain = 5

01:07:42 - MC IN2 = 8dB, CARM gain = 6.295  (AO went up to 6dB, then +1dB steps to both simultaneously using  ezcastep C1:LSC-CARM_GAIN 1dB C1:IOO-MC_AO_GAIN 1)

01:08:57 - MC IN2 = 10dB, CARM gain = 7.92447

01:10:08 - MC IN2 = 12dB, CARM gain = 9.97631

lockloss when trying to add 1 more dB to both.

01:41:36 - MC IN2 = 12dB, CARM gain = 9.97

lockloss when just MC IN2 up by 1dB, left CARM gain alone.


Other notes:

The 60Hz noise in TRY is back.  Since I thought I remembered someone suggesting that it was leakage light from the exit sign, Koji went in and wrapped the end table in foil, however the lines are still present.

  9955   Thu May 15 01:42:07 2014 ranaUpdateCDS/frames space cleared up, daqd stabilized

 Script seems to be working now:

nodus:~>df -h | grep frames

fb:/frames              13T    12T   931G    93%    /frames

  9954   Wed May 14 17:36:32 2014 ericqUpdateCDSNew netgpib scripts for SR785

 I have redone the SPSR785 (spectrum measurement) and TFSR785 (TF measurement) commands in scripts/general/netgpibdata. This was mostly motivated by my frustration with typing out either a ton of command line arguments, or rooting around in the script itself; I'd rather just have a static file where I define the measurement, and can keep track of easier. 

They currently take one argument: a parameter file where all the measurement details are specified. (i.e. IP address, frequencies, etc.) There are a few template files in the same directory that they use as default. (Such as TFSR785template.yml)

If you call the functions with the option '--template', it will copy a template file into your working directory for you to modify as you wish. "SPSR785 -h" gives you some information as well (currently minimal, but I'll be adding more)

In the parameter file, you can also ask for the data to be plotted (and saved as pdf) when the measurement is finished. In SPSR785, and soon TFSR785, you can specify a directory where the script will look for reference traces to plot along with the results, presuming they were taken with the same measurement parameters and have the same filename stem. 

I've tested both on Pianosa, and they seem to work as expected. 

Todo:

  • Add support for modifying some parameters at the command line 
  • Extend to the Agilent analyzer
  • Maybe the analyzer settings written to the output file should be verified by GPIB query, instead of writing out the intended settings. (I've never seen them go wrong, though)
  • Make sure that the analyzer has PSD units off when taking a TF. (Thought I could use resetSR785 for this, but there's some funkiness happening with that script currently.)
  • Possibly unify into one script that sees what kind of analyzer you're requesting, and then passes of to the device/measurement type specific script, so we don't have to remember many commands. 

Comments, criticism, and requests are very welcome. 

(P.S. all the random measurement files and plots that were in netgpibdata are now in netgpibdata/junk. I feel like this isn't really a good place to be keeping data. Old versions of the scripts I changed are in netgpibdata/oldScripts)

  9953   Wed May 14 14:39:31 2014 JenneUpdateSUSNew buttons on IFO_ALIGN screen

It's starting to get a little crowded, but I modified the IFO_ALIGN screen to have new buttons to show the aligned / misaligned state of each optic.  Koji made a good point, and I left the old restore script functional so that if the slider is moved significantly, we can always go back gently to the burt restored value.  I have removed the old misalign function though, since we shouldn't ever be using that again.

Screenshot-Untitled_Window.png

  9952   Wed May 14 10:04:06 2014 SteveSummarySUS oplev laser and temperature

 ETMX oplev qpd gain has to be increased.

 

Attachment 1: TempRules.png
TempRules.png
  9951   Wed May 14 02:37:58 2014 JenneUpdateLSCNo big locking news

Tonight was mostly cleaning up some scripts, including the re-writing the restore and align scripts for the optics. 

The new script is in the same folder as the old one (/opt/rtcds/caltech/c1/medm/MISC/ifoalign/NewAlignSoft.py), but is not yet called from the align screen.  However, I'm using it in the carm up and down scripts, and it works nicely for the PRM.  I need to check that the offset value is okay for all the other optics (i.e. are they getting misaligned enough?), but then I'll have the new script called from the screen.  The new script, per Rana's suggestion, does not touch the bias sliders.  Rather, it puts an offset in the pitch filter banks in the coil driver output matrix-of-filter-banks.  Then the misalign routine turns the offset on, and the restore routine turns the offset off.  This way we have a nice ramp time, but don't have to do the weird calculation of number of steps to take as is done in the current script.  Also, the "save" functionality will be obsolete, since we're never touching the bias sliders except for actual alignment needs.

I'm not sure what changed, other than the HEPA being on lower, but the Xarm ALS was much better behaved tonight.  I was able to hang out around arm powers of ~1 for as long as I wanted. 

I didn't try to hand over to digital REFLDC, but I was trying a few times to engage the AO path.  With the CM board set to Plus, I hear hooting when MC IN2 is about 4dB.  With the CM board set to Minus, I didn't hear hooting, but I lost lock when I went from 13dB to 14dB. 

Also, I put the cables for the SR785 back to the "A" set of test points and excitation, so that I could take a closed loop transfer function.  However, I don't know where the latest working scripts to make a remote measurement are, so maybe we can take some loop measurements tomorrow.

The carm_cm_up script is good (for tonight) up to the prompt "Press enter to indicate that it is okay to turn on MC2 LSC FM8".  There are "read"s every step of the way, so it goes nice and slow, but it'll do everything for you except any last tweaks of the PRM alignment after the PRMI is locked.

  9950   Tue May 13 22:55:57 2014 JenneSummarySUSETMX oplev: cleanup

I believe that the Xend aux laser was turned off earlier today, for Steve's work swapping out the oplev.  When I went down there, the red "off" LED was illuminated, and the LCD screen was showing something.  I pushed the green "on" button, and I immediately got green.

Also, I saw that the 24Hz roll mode was very rung up on ETMX.  I looked at the FM5 "bounce roll" filter, and it had some old values, 12Hz and 18Hz for the resonant gains.  All other optics have the proper 16Hz and 24Hz frequencies.  I copied the BS oplev bounce roll filter over to ETMX pit and yaw (both were wrong), and loaded them in.  The mode is starting to ring down.

  9949   Tue May 13 17:45:21 2014 ranaUpdateCDS/frames space cleared up, daqd stabilized

 

 Late last night we were getting some problems with DAQD again. Turned out to be /frames getting full again.

I deleted a bunch of old frame files by hand around 3AM to be able to keep locking quickly and then also ran the wiper script (target/fb/wiper.pl).

controls@pianosa|fb> df -h; date

Filesystem            Size  Used Avail Use% Mounted on

/dev/sda1             440G  9.7G  408G   3% /

none                  7.9G  288K  7.9G   1% /dev

none                  7.9G  464K  7.9G   1% /dev/shm

none                  7.9G  144K  7.9G   1% /var/run

none                  7.9G     0  7.9G   0% /var/lock

none                  7.9G     0  7.9G   0% /lib/init/rw

none                  440G  9.7G  408G   3% /var/lib/ureadahead/debugfs

linux1:/home/cds      1.8T  1.4T  325G  82% /cvs/cds

linux1:/ligo           71G   18G   50G  27% /ligo

linux1:/home/cds/rtcds

                      1.8T  1.4T  325G  82% /opt/rtcds

fb:/frames        13T   12T  559G  96% /frames

linux1:/home/cds/caltech/users

                      1.8T  1.4T  325G  82% /users

Tue May 13 17:35:00 PDT 2014

Looking through the directories by hand it seems that the issue may be due to our FB MXstream instabilities. The wiper looks at the disk usage and tries to delete just enough files to keep us below 95% full for the next 24 hours. If, however, some of the channels are not being written because some front ends are not writing their DAQ channels to frames, then it will misestimate the disk size. In particular, if its currently writing small frames and then we restart the mxstream and the per frame file size goes back up to 80 MB, it can make the disk full.

For now, I have modified the wiper.pl script to try to stay below 93%. As you can see by the above output of 'df', it is already above 96% and it still has files to write until the next run of wiper.pl 7 hours from now at. at 6 AM.

IF we assume that its writing a 75MB file every 16 seconds, then it would write 405 GB of frames every day. There is 559 GB free right now so we are OK for now. With 405 GB of usage per day, we have a lookback of ~12TB/405GB ~ 29 days (ignoring the trend files).

  9948   Tue May 13 17:31:32 2014 JenneSummarySUSETMX oplev laser replaced: New oplev gains set

I took loop measurements of ETMX pit and yaw, and set the upper UGF to be ~6Hz for both.  This required a pitch gain of 25, and a yaw gain of 16.

The spectra look similar to what they were before Steve did the swap.

OLerr_13May2014.pdf

OLfb_13May2014.pdf

  9947   Tue May 13 17:03:05 2014 SteveSummarySUSETMX oplev laser replaced

Quote:

Quote:

For some reason or another, I decided that we should see if the optical lever servos were injecting too much noise into the test masses. The ITMs are much worse than the ETMs and I am afeared that they might be making the main noise for our arms in the 20-40 Hz region. Jenne is checking up on these feedback loops to see what's up.

To estimate the actuator gains of the mirrors, I turned on 1 count drives from LSC/CAL oscillators into the LSC drives of each test mass at the frequencies shown in the plot with the resulting peaks showing up in in POX/Y with the single arm locks in red. I will leave these going permanently, but with 0.1 count ampltiudes; we need to make it so in the scripts.

 I'm in the process of filling this table

OPLEV

SERVO     

300 ^

2:0

BR ELP RLP BOOST RES

GAIN

QPD

COUNTS

 

QPD

mW

QPD

beam

OD

HE/NE

output

mW

%

back

on QPD

                                
ETMY PIT  FM1  FM5    55      -30 8,200 0.2   3.3  
        YAW  FM1  FM5    55      -28          
ETMX PIT  FM1  FM5  35        4.4 900 0.2   1.7  
       YAW  FM1  FM5  35       2.1  1,750  0.33    2.8  
ITMY PIT  FM1  FM5        3.3   52 14,400 0.4   9.5  
      YAW  FM1  FM5        3.3  -46          
ITMX PIT  FM1  FM5  50      3.3   30 7,400 0.17   2.8  
       YAW  FM1  FM5  50      3.3  -20          
BS  PIT  FM1  FM5  35      3.3   9 2,800 0.05   2.8  
    YAW  FM1  FM5  35      3.3  -9          
PRM  PIT FM1  FM5  55   FM7  3.3  7 3200 0.06   2.8  
       YAW FM1  FM5  55   FM7  3.3  -5          
SRM  PIT  FM1            -20       9.5  
       YAW  FM1            20          

 I should replace ETMX He/Ne laser

 

  9946   Tue May 13 13:27:58 2014 SteveSummarySUSOptical Lever Servos setting table

Quote:

For some reason or another, I decided that we should see if the optical lever servos were injecting too much noise into the test masses. The ITMs are much worse than the ETMs and I am afeared that they might be making the main noise for our arms in the 20-40 Hz region. Jenne is checking up on these feedback loops to see what's up.

To estimate the actuator gains of the mirrors, I turned on 1 count drives from LSC/CAL oscillators into the LSC drives of each test mass at the frequencies shown in the plot with the resulting peaks showing up in in POX/Y with the single arm locks in red. I will leave these going permanently, but with 0.1 count ampltiudes; we need to make it so in the scripts.

 I'm in the process of filling this table

OPLEV

SERVO     

300 ^

2:0

BR

16,24

Hz

ELP RLP BOOST RES

GAIN

QPD

COUNTS

 

QPD

mW

QPD

beam

OD

HE/NE

output

mW

%

back

on QPD

                                
ETMY PIT  FM1  FM5    55      -30 8,200 0.2   3.3  
        YAW  FM1  FM5    55      -28          
ETMX PIT  FM1  FM5  35        4.4 900 0.2   1.7  
       YAW  FM1  FM5  35        2.1          
ITMY PIT  FM1  FM5        3.3   52 14,400 0.4   9.5  
      YAW  FM1  FM5        3.3  -46          
ITMX PIT  FM1  FM5  50      3.3   30 7,400 0.17   2.8  
       YAW  FM1  FM5  50      3.3  -20          
BS  PIT  FM1  FM5  35      3.3   9 2,800 0.05   2.8  
    YAW  FM1  FM5  35      3.3  -9          
PRM  PIT FM1  FM5  55   FM7  3.3  7 3200 0.06   2.8  
       YAW FM1  FM5  55   FM7  3.3  -5          
SRM  PIT  FM1            -20       9.5  
       YAW  FM1            20          

 I should replace ETMX He/Ne laser

Attachment 1: OLsums.png
OLsums.png
  9945   Tue May 13 03:30:10 2014 JenneUpdateLSCLocking activities - no progress

[Jenne, Rana, EricQ]

We tried a few times to engage the AO path while holding CARM on sqrtTrans and DARM on ALS, but failed every time.  Since we cannot stably hold lock at arm powers of 1, even though we were able to do so early last week, we think that we have a problem (obviously).  One noticeable thing is that while held with ALS, the Xarm transmission fluctuates almost full power.

As we were seeing late last week, the Xarm IR transmission while held with ALS was fluctuating wildly, whether we were locked on individual arms or on CARM/DARM. 

Tonight we took some out of loop spectra, with different HEPA settings, to see how that affected things.  It looks like HEPA at the nominal ~20% is okay, but anything higher than that starts to affect the Xarm beatnote sensing, while it mostly leaves the Yarm beatnote sensing okay.  Perhaps this is something that isn't tightened enough in the Xarm path, or something on a skinny floppy mount that needs to be more secure.

I am still a little confused though, why we don't see large power fluctuations in *both* arms while using ALS to control CARM/DARM.  Why are we not seeing this Xarm noise being fed back into the Yarm, through either the ETM via DARM, or common stuff via CARM actuating on MC2.

Note that the change at high frequency is because I switched from using non-DQ channels to DQ channels, so that's not anything to pay attention to.  The noise reduction we see is below about 20Hz.

ALS_outofloop.pdf


Rana pointed out that our POPDC level was very small.  We don't have screens for them, but the DC PDs have the same analog whitening as the RF PD signals do.  I changed C1:LSC-POPDC_WhiteGain.VAL from 0 to 11.  Now our POPDC while locked on sidebands is about 8,000 counts.

We also swapped the cables between the SR785 and the CM board around.  Now channel 2 of the 785 goes to TP2A, channel 1 goes to TP2B, and the source goes to EXCB.  This is so that we can break the AO loop with the disable switch just after the slow/fast split, and look at the transfer function before we close the loop. 

  9944   Tue May 13 00:46:58 2014 ranaHowToPSLPMC relocking

The PMC runs out of range sometimes due to the daily temperature swing. The voltage swings up after sunset and then starts to swing down before sunrise. So when you relock the PMC at the beginning of the locking night, the mnemonic from the PMC is:

Sun Go Low, Lock Me Voltage Low.

  9943   Mon May 12 22:52:59 2014 JenneSummarySUSOptical Lever Filters are all different

We need to go back and have a look at all of our optical lever control filters, and make sure they make sense. 

In particular, we should have a look at the ITMs, since they have a huge amount of motion around 10Hz. 

Notes:  ETMX shouldn't have that lower notch.  The bounce mode Qs should be lowered.

OpLevFilters.pdf

  9942   Mon May 12 22:42:19 2014 ranaSummarySUSOptical Lever QPD Sum trends: they're almost all too weak

For some reason or another, I decided that we should see if the optical lever servos were injecting too much noise into the test masses. The ITMs are much worse than the ETMs and I am afeared that they might be making the main noise for our arms in the 20-40 Hz region. Jenne is checking up on these feedback loops to see what's up.

To estimate the actuator gains of the mirrors, I turned on 1 count drives from LSC/CAL oscillators into the LSC drives of each test mass at the frequencies shown in the plot with the resulting peaks showing up in in POX/Y with the single arm locks in red. I will leave these going permanently, but with 0.1 count ampltiudes; we need to make it so in the scripts.

Attachment 1: OL-FB.png
OL-FB.png
Attachment 2: arms_140512.pdf
arms_140512.pdf
  9941   Mon May 12 14:42:25 2014 steveUpdateendtable upgradeoptical table enclosure wall proposal and table at ETMY

Quote:

 Sanwiched wall as shown: 1" clear acrylic, 2 layers of 0.004" thick "window tint", 1 layer of  0.007" thermashield  and  0.125" yellow acrylic

Visibility: 70 %,    Transmission of 1064 nm  2-3 % at 0-50 degrees incident,   power density  ~ 0.7 W/mm2

Max power 100 mW

        

More details about this east end  " acrylic + "   enclosure ( optical table cover ) can be found elog entry 6210, 7194 and  7106

 

Window tinted layer transmission plot is below.

 

We have a film which may meet your requirements and the values are shown below:

         

               Wavelength (nm)              Transmission                     Reflectance(front)            Reflectance (back)

 

                        1060                               .0772                               .604                                     .759

 

                        1070                               .0723                               .615                                     .772

 

These values are taken from the LBNL Optics 6 program and if you have access to that program, the NFRC ID for the film is 202.   If you do not have access to the program, I have a captured the graph which may be of some help.  I apologize for the appearance of the graph but someone at LBNL decided it would look better with a dark gray background – the yellow is the transmission curve, the blue is the reflectance (front) and the green is the reflectance (back).

 

 

The film is referred to as “Hilite 70” and has a 72% visible light transmission.  These results were obtained with the film mounted on 1/8” clear glass.

 

I

 

 

 

 

 

Saint-Gobain Performance Plastics

www.solargard.com

 

Please consider your environmental responsibility before printing this email.

 

 

 

Attachment 1: Hilite70.jpg
Hilite70.jpg
Attachment 2: ETMY-ISCT.jpg
ETMY-ISCT.jpg
Attachment 3: RoscoThermashield.pdf
RoscoThermashield.pdf
ELOG V3.1.3-